OpenStudy (anonymous):

What is the slope of the line passing through the points (6, -4) and (8, -7)? is this -3/2?

OpenStudy (anonymous):

We plug it in the slope formula: \(m = \dfrac{y_2-y_1}{x_2-x_1}\) \(m = \dfrac{-7+4}{8-6}\) \(m = -\dfrac{3}{2}\) Nice work! \(\Huge\color{lime} \checkmark\)